In a steam engine the steam at the beginning of the expansion process is

at 7 bar, dryness fraction 0.98 and expansion follows the law pv1.1 = constant, down to a pressure

of 0.34 bar. Calculate per kg of steam :

(i) The work done during expansion ;

(ii) The heat flow to or from the cylinder walls during the expansion.


1
Expert's answer
2021-08-29T16:55:48-0400

Initial pressure of steam,

"p_{1} =7 \;bar = 7 × 10^{5} \; N/m^{2}"

Dryness fraction,

"x_{1} = 0.98"

Law of expansion,

"pv^{1.1} = constant"

Final pressure of steam,

"p_{2} =0.34 \; bar = 0.34 \times 10^{5} \; N/m^{2}"

At 7 bar :

"v_{g} = 0.273 \; m^{3}/kg \\

v_{1}=x_{1}v_{g}= 0.98 \times 0.273 = 0.267 \; m^{3}/kg \\

p_{1}v_{1}^{n}=p_{2}v_{2}^{n} \\

\frac{v_{2}}{v_{1}}=\left ( \frac{p_{1}}{p_{2}} \right )^{\frac{1}{n}} \\

\frac{u_{2}}{0.267}=\left ( \frac{7}{0.34} \right )^{\frac{1}{1.1}}"

or

"v_{2}= 0.267\left ( \frac{7}{0.34} \right )^{\frac{1}{1.1}}= 4.174 \; m^{3}/kg"

(i) Work done by the steam during the process :

"W=\frac{p_{1}v_{1}-p_{2}v_{2}}{n-1} \\

=\frac{7\times 10^{5}\times 0.267-0.34\times 10^{5}\times 4.174}{\left ( 1.1-1 \right )} \\

=\frac{10^{5}}{0.1}(1.869 - 1.419) =10^{5}\times 4.5 \; Nm/kg"

Work done "= \frac{10^{5}\times 4.5}{10^{3}} =450\; kJ/kg."

At 0.34 bar : "v_{g}= 4.65\; m^{3}/kg" , therefore, steam is wet at state 2 "(since \;v_{2} < v_{g})" .

Now, "v_{2}=x_{2}v_{g}" , where "x_{2} =" dryness fraction at pressure "p_{2} \; (0.34 \;bar)"

"4.174 = x_{2}\times 4.65"

or

"x_{2} = \frac{4.174}{4.65}= 0.897"

The expansion is shown on a p-v diagram in Fig., the area under 1-2 represents the work done per kg of steam.



(ii) Heat transferred :

Internal energy of steam at initial state 1 per kg,

"u_{1}= (1 - x_{1})u_{f} + x_{1}u_{g} \\

= (1 - 0.98) 696 + 0.98 \times 2573 = 2535.46 \; kJ/kg"

Internal energy of steam at final state 2 per kg,

"u_{2}=\left ( 1-x_{2} \right )u_{f}+x_{2}u_{g} \\

= (1 - 0.897) 302 + 0.897 \times 2472 = 2248.49 \; kJ/kg"

Using the non-flow energy equation,

"Q=\left ( u_{2} -u_{1}\right )+W \\

= (2248.49 - 2535.46) + 450 = 163.03 \; kJ/kg"

Heat supplied = 163.03 kJ/kg.